# Are foundation models for computer vision good conformal predictors?

### v1

@@ Abstract
Recent advances in self-supervision and constrastive learning have brought the performance of foundation models to unprecedented levels in a variety of tasks. Fueled by this progress, these models are becoming the prevailing approach for a wide array of real-world vision problems, including risk-sensitive and high-stakes applications. However, ensuring safe deployment in these scenarios requires a more comprehensive understanding of their uncertainty modeling capabilities, which has been barely explored. In this work, we delve into the behavior of vision and vision-language foundation models under Conformal Prediction (CP), a statistical framework that provides theoretical guarantees of marginal coverage of the true class. Across extensive experiments including popular vision classification benchmarks, well-known foundation vision models, and three CP methods, our findings reveal that foundation models are well-suited for conformalization procedures, particularly those integrating Vision Transformers. Furthermore, we show that calibrating the confidence predictions of these models leads to efficiency degradation of the conformal set on adaptive CP methods. In contrast, few-shot adaptation to downstream tasks generally enhances conformal scores, where we identify Adapters as a better conformable alternative compared to Prompt Learning strategies. Our empirical study identifies APS as particularly promising in the context of vision foundation models, as it does not violate the marginal coverage property across multiple challenging, yet realistic scenarios.
@@ 1 Introduction
Large-scale pre-trained vision foundation models, such as DINOv2 [33], as well as those integrating text, such as CLIP [38], are driving a new learning paradigm in machine learning, achieving unprecedented results on a broad spectrum of tasks. Nevertheless, despite their desirable zero-shot and generalization capabilities, recent evidence has pointed out to the existence of bias and factual errors in these models [49], which transcend the field of computer vision [44, 4]. For example, the original CLIP paper [38] demonstrated gender and race biases in certain zero-shot tasks, whereas [49] identified that CLIP-based models are not always better calibrated than other arguably simpler ImageNet models. Furthermore, [30] recently showed that adapted models magnify the miscalibration issue compared to the zero-shot setting, yielding overconfident predictions. These problems underscore widespread societal concerns surrounding the reliable deployment and use of foundation models in sensitive contexts, such as decision-making processes in critical scenarios, e.g., healthcare, autonomous driving or security applications.
@@
A popular solution to quantify the uncertainty present in the predictions of deep models is calibration. In this setting, the proposed strategies aim at reducing the discrepancies between model predictions and the actual correctness probability. Temperature Scaling (TS) [13], a simple variant of Platts Scaling [36], provides a simple post-processing approach to adjust the softmax probability scores of the trained models. Other line of methods have proposed training objectives to enforce the model to produce less confident scores, either in the predictions space [37, 35], logits [27, 26], or modifying the ground truth labels [28, 29]. Nevertheless, confidence calibration does not offer theoretical guarantees for coverage of the true class.
@@
In contrast, Conformal Prediction (CP) [51], an alternative strategy to quantify the uncertainty [41, 40, 1], is a statistical framework which offers several advantages over calibration methods. First, unlike most methods in calibration, CP works directly on the model predictions, offering an appealing solution for black-box models. Second, instead of simply modeling the correctness of the predicted probability, CP produces a set of predictions, including the most likely classes, which can be of much interest in certain problems. Last, CP methods have theoretical guarantees for the marginal coverage of the true class within the predicted set, under several assumptions.
@@
Due to these properties, CP is gaining attention within the community to conformalize the predictions of deep models [2, 1, 9, 42]. Nevertheless, albeit the efforts to study CP in large language foundation models [12], its impact on vision foundation models has been unexplored, besides the significant implications it may have on a variety of vision problems. The aim of this study is to shed light and provide important insights into this direction. To achieve this, we conducted an extensive empirical analysis of the performance of three common CP methods on ten popular vision foundation models across three vision datasets. Our extensive experiments further explore common situations encountered in practice, assessing the impact on CP: under distributional drifts, after confidence calibration and in few-shot adaptation to novel downstream tasks. Our key observations are:
@@
    Across all the experiments, Adaptive Prediction Sets (APS) is the best Conformal Prediction approach in terms of empirical coverage, while Regularized Adaptive Prediction Sets (RAPS) presents the best alternative from a conformal set size standpoint.
@@
    Adapters are generally better candidates for CP than Prompt Learning methods, in both in-distribution and out-of-distribution scenarios.
@@ 2 Related Work
Foundation models for computer vision.  The landscape of foundation models has rapidly evolved in the last years. Traditionally, pretrained convolutional networks based on ResNet architectures [15] were the main models used by the community. However, recently driven by the unprecedented advances in language models, such as GPT [6] or LLaMA [48], as well as the vast availability of image data online, there are groundbreaking advances in unimodal [7, 33, 22] and multimodal [38] foundation models for vision tasks based mainly on vision transformer architectures. These large pre-trained models aim to generalize across a broad range of visual tasks by pre-training on massive, diverse image datasets, exhibiting strong zero-shot and generalization capabilities to new tasks. For example, vision foundation models, such as DINO [7, 33], rely on self-supervised learning strategies on large datasets, which results in an excellent semantic understanding of the visual content. On the other hand, CLIP [38] bridges the gap between language and vision modalities through contrastive learning, effectively allowing the model to understand images in the context of natural language prompts and enabling zero-shot capabilities.
@@
Uncertainty estimation in deep models.  Quantifying the uncertainty of the predictions of deep networks has recently garnered considerable interest. From a calibration standpoint, popular strategies include post-hoc approaches [13, 10, 20], which map the logits or softmax predictions to smoother distributions, and explicit learning objectives [26, 27, 29, 28, 16, 31], which are integrated into the loss function. Nevertheless, a main limitation of calibration methods is that they lack theoretical guarantees of model performance. In contrast, CP has recently emerged as a promising alternative, which provides marginal coverage guarantees over unseen test samples [52, 43]. Specifically, CP resorts to a non-conformity score function (i.e., a measure of how “different” a particular data point is compared to a CP calibration dataset) to produce a finite prediction set, which is guaranteed to contain the true label with a specified confidence level. A central objective of the CP literature has been to improve either the set efficiency (i.e., smaller set sizes) or the class conditional coverage. For this purpose, several non-conformity scores have been presented in the literature [47, 40, 1, 41, 11, 9, 46], with [40, 1, 41] being popular methods widely studied. A straighforward solution consists on directly using the raw class softmax predictions to generate the prediction sets [41]. Adaptive Prediction Sets (APS) [40] provides an adaptive version, by computing the non-conformity scores by accumulating sorted softmax probabilities in descending order. To further improve the efficieny, RAPS [1] introduces an explicit regularization term, which penalizes non-conformity scores for unlikely classes.
@@
However, a main limitation of existing evaluations is the focus on more traditional models, usually trained on a data collection that falls within the distribution of calibration and test data points. Thus, given the rising popularity of foundation models in strategic domains, quantifying the uncertainty of their predictions is paramount. However, whereas uncertainty quantification from a calibration perspective has been scarcely studied [30, 55, 50], its exploration under CP is, to the best of our knowledge, overlooked.
@@ 3.2 Non-conformity scores
Least Ambiguous Classifier.  LAC [41] aims to construct the smallest possible set under the assumption that the output is correct. Intuitively, it can be interpreted as a thresholding of the output probabilities for each category. Thus, the non-conformity score can be constructed as:
@@
where $\rho(\mathbf{x},y)$ is the accumulated confidence of the categories more likely than the evaluated label $y$, i.e., $\rho(\mathbf{x},y)=\sum_{k^{\prime}\in\mathcal{Y^{\prime}}(\mathbf{x},y)}x_{k=k^{\prime}}$, with $\mathcal{Y^{\prime}}(\mathbf{x},y)=\{k|x_{k}>x_{k=y}\}$. Adaptive methods usually include $u\in\{0,1\}$, as a random variable to break ties to archive exact marginal coverage.
@@ 4.1 Experimental Setup
Models: We employ a total of ten foundation models: two DINO [7] (DINO-S and DINO-B), four DINOv2 [33] (DINOv2-S, DINOv2-B, DINOv2-L, and DINOv2-G), three VICReg [3] (based on ResNet-50, ResNet-50x2, and ResNet-200x2), and one vision-language model (CLIP [38] based on ViT-B). Our analysis is conducted on three popular vision datasets: CIFAR-10 [23], CIFAR-100 [23], and Imagenet [8], including its versions integrating domain shifts [17, 18, 53, 39]. Each dataset is split into two sets: one for training the linear probing head, and one for the conformal experiments. The latter is then split into one calibration set to tune the CP method, and one test set for evaluation.
Metrics: We resort to conformal set size (a.k.a efficiency), minimum class-conditional coverage (MCCC), coverage gap (CovGap) and marginal coverage to benchmark the different methods. More details about metrics and implementation details can be found in Appendix A and Appendix B.
Figure 1: Relationship between the linear probing model accuracy and conformal set size (top) and the coverage gap (bottom) across different tasks of increasing complexity. From left to right: CIFAR-10, CIFAR-100 (middle) and ImageNet.
Adaptation to target tasks.  The foundation models used in this study have been pre-trained using different strategies. Nevertheless, they need to be adapted for novel tasks, as their pre-trained versions do no accommodate classification tasks, i.e., there is no classification head. To do this, foundation models are frozen and a Linear Probing head (one linear layer followed by a softmax activation function) is trained on each dataset by optimizing a cross-entropy loss.
@@ 4.2 Results
To gain insights into the factors influencing the efficacy of CP approaches in vision foundation models, we design four experiments. First, we start by exploring the impact of CP in standard scenarios, where a large calibration set is available to conformalize the predictions of different models. Then, we challenge the status-quo of CP, and alter the conformal sets to accommodate real-world scenarios, by including domain shifts. Furthermore, the notion of confidence calibration being very linked to CP, we explore the impact of calibrating a model on the performance of CP. Last, we explore the impact of CP when adapting a very popular VLM, i.e., CLIP, to novel tasks.
@@ 4.2.1 Performance in the General Setting
First, we study the performance of ten vision foundation models paired with CP methods under the standard setting, and on the three datasets, which presents ideal conditions: a sufficiently large calibration set, and absence of distributional drifts between calibration and test sets. Our goal is to determine whether we can prescribe a winner solution in this scenario, and which factors can help its identification.
@@
Fig. 1 depicts the relationship between the linear probing performance of each model and the conformal set size (top), and coverage gap (bottom). The initial observation highlights a clear trend: models with higher performance tend to produce smaller prediction sets, regardless of the conformal method used. Nevertheless, while set efficiency (i.e. size) is typically considered as a sufficient condition in most prior literature in conformal prediction [1], our analysis revealed that the relation with the accuracy does not consistently hold when examining other metrics. In particular, Fig. 1 (bottom) exposes that different CP methods yield mixed results for the coverage gap, which do not correlate directly with accuracy across all CP approaches. While APS appears to be almost unaffected by the model performance, RAPS clearly benefits from a strong performance, and LAC is negatively affected by more accurate models as the dataset becomes more complex.
@@
In terms of comparison between the CP methods, if we consider the set size, APS is clearly outperformed by the other approaches, particularly on ImageNet, whereas LAC provides the smallest prediction sizes, closely followed by RAPS. Indeed, RAPS is specifically designed to reduce the conformal set size of APS. However, the coverage gap results indicate that this comes at the cost of increasing the range of the class-conditional coverage. Below, we analyze the underlying causes that may explain this behaviour.
@@
RAPS class-conditional coverage, and therefore coverage gap, are more sensitive to the accuracy of the model. Let us assume we have two models, $M_{1}$ and $M_{2}$, in a multi-class classification problem, whose accuracies are $Acc_{1}$ and $Acc_{2}$, respectively. For each class $y\in\mathcal{Y}$, we refer to $C_{M_{i}}(y)=\mathbb{P}(Y=y|Y\in\mathcal{S}_{M_{i}}(X))$ as the class-conditional coverage for $y$ under model $M_{i}$, which measures the probability that predictions include the true label when the true label is $y$. Furthermore, let $\delta_{M_{1}}=\min_{y\in\mathcal{Y}}C_{M_{1}}(y)$ and $\delta_{M_{2}}=\min_{y\in\mathcal{Y}}C_{M_{2}}(y)$ denote the minimum class-conditional coverage achieved by each model. Under this scenario, we argue that due to the penalty in RAPS, if $Acc_{1}<Acc_{2}$, then $\delta_{M_{1}}<\delta_{M_{2}}$ when using RAPS as conformal prediction method. In particular, $M_{1}$ (with lower accuracy) needs to expand its prediction sets for certain classes in order to meet the marginal coverage target 1-$\alpha$. Nevertheless, the penalty term encourages small prediction sets, limiting an excessive number of classes. Thus, for some difficult classes, model $M_{1}$ may still potentially fail to meet the target coverage, as the enforced penalty discourages overly large sets. This ultimately results in lower coverage rates for those specific classes compared to $M_{2}$. In contrast, since APS does not include any regularization term that encourages small set sizes, it will compensate for lower performing models by increasing its set sizes, ultimately attaining higher class-conditional coverages.
@@
To confirm this hypothesis, we perform the following experiment across all foundation models. First, we identify the class with the lowest minimum class-conditional coverage obtained by RAPS for each model $m$, which we refer to as $k^{R}_{m}$, and find its corresponding set size (both values are represented by pink circles in Fig. 2). Then, we identify the class-conditional coverage and set size of the class $k^{R}_{m}$ provided by APS across all models, whose values are shown as blue circles in both plots of Fig. 2. Last, we also include the minimum class-conditional coverage obtained by APS across each method, depicted with a cross. Note that the minimum class-conditional coverage from APS does not necessarily correspond to the coverage of class $k^{R}_{m}$. Upon close examination of these results, we observe that, indeed, for models presenting lower accuracy (i.e., those with smaller circles), the gap between the class-conditional coverage for the worst class in RAPS and the same class in APS is consistently larger than for more accurate models (larger circles).
This empirical evidence shows that, because RAPS is constrained from expanding its predictions set size, the minimum class-conditional coverage is substantially reduced. This effect is more pronounced in less accurate models, where the true class may rank far away from the maximum allowable set size in the softmax predictions. In contrast, APS tolerates better worse models by increasing the set size, which ultimately reduces the efficiency of the sets but yields better class-conditional coverages. Last, LAC presents structural differences with RAPS and APS, as it lacks an adaptive mechanism and instead relies on a uniform fixed threshold. Thus, LAC may yield inconsistent coverage rates across classes, resulting in high variability in the class-conditional coverage and thus in the coverage gap.
@@
Table 1: Results on ImageNet obtained by CLIP (based on ViT-B) and a ViT-B trained in a supervised manner on ImageNet, across accuracy and min class-conditional coverage (MCCC).
Are foundation models better conformable networks? Following this analysis, we are also interested in determining whether a network pre-trained following a more traditional approach (i.e., standard supervised fine-tuning) offers similar conformal capabilities to self-supervised ones. In particular, we select a ViT-B pre-trained on ImageNet, which is the same architecture as the visual encoder of the CLIP model, so the potential differences may come from the different learning strategies, as well as the amount of employed data. The results from this analysis (Table 1) reveal that, despite obtaining lower classification accuracy when using LP on the $\text{ViT}_{\text{CLIP}}$ model, CP methods typically yield to better performance than in $\text{ViT}_{\text{ImageNet}}$. These differences are particularly important under the APS approach, where the set size is significantly degraded on $\text{ViT}_{\text{ImageNet}}$. Moreover, the class-conditional coverage is also substantially affected, particularly for APS, with nearly 6% decrease.
@@
Figure 4: Evaluation under domain-shift. Set size($\downarrow$), coverage($\uparrow$), and MCCC($\uparrow$) across three CP methods and three foundation models. ImageNet versions are sorted based on their OOD performance in [45].
To further delve into these differences, we compute, for each test sample, the difference between the conformal set size when applied to $\text{ViT}_{\text{ImageNet}}$ and $\text{ViT}_{\text{CLIP}}$ models, whose distribution is depicted in Fig. 3. These values confirm that set size differences are not derived from a small set of isolated outliers, but a considerably large group of samples which see their conformal set increase when using the ViT trained in a supervised manner. These results suggest that the strategies used to train foundation models tend to yield better CP metrics, resulting in conformalized models that can be deployed more safely on critical scenarios.

### v2

@@ Abstract
Recent advances in self-supervision and contrastive learning have brought the performance of foundation models to unprecedented levels in a variety of tasks. Fueled by this progress, these models are becoming the prevailing approach for a wide array of real-world vision problems, including risk-sensitive and high-stakes applications. However, ensuring safe deployment in these scenarios requires a more comprehensive understanding of their uncertainty modeling capabilities, which has been barely explored. In this work, we delve into the behaviour of vision and vision-language foundation models under Conformal Prediction (CP), a statistical framework that provides theoretical guarantees of marginal coverage of the true class. Across extensive experiments including popular vision classification benchmarks, well-known foundation vision models, and three CP methods, our findings reveal that foundation models are well-suited for conformalization procedures, particularly those integrating Vision Transformers. We also show that calibrating the confidence predictions of these models, a popular strategy to improve their uncertainty quantification, actually leads to efficiency degradation of the conformal set on adaptive CP methods. Furthermore, few-shot adaptation of Vision-Language Models (VLMs) to downstream tasks, whose popularity is surging, enhances conformal scores compared to zero-shot predictions. Last, our empirical study exposes APS as particularly promising in the context of vision foundation models, as it does not violate the marginal coverage guarantees across multiple challenging, yet realistic scenarios.
@@ 1 Introduction
Large-scale pre-trained vision foundation models, such as DINOv2 [40], as well as those integrating text, such as CLIP [46], are driving a new learning paradigm in machine learning, achieving unprecedented results on a broad spectrum of tasks. Despite their desirable zero-shot and generalization capabilities, recent evidence has pointed out to the existence of bias and factual errors in these models [58], which transcend the field of computer vision [52, 4]. For example, the original CLIP paper [46] demonstrated gender and race biases in certain zero-shot tasks, whereas [58] identified that CLIP-based models are not always better calibrated than other arguably simpler ImageNet models. Furthermore, [36] recently showed that adapted models magnify the miscalibration issue compared to the zero-shot setting, yielding overconfident predictions. These problems underscore widespread societal concerns surrounding the reliable deployment and use of foundation models in sensitive contexts, such as decision-making processes in critical scenarios, e.g., healthcare or security applications.
@@
A popular solution to quantify the uncertainty present in the predictions of deep models is calibration. In this setting, the proposed strategies aim at reducing the discrepancies between model predictions and the actual correctness probability. Temperature Scaling (TS) [16], a simple variant of Platts Scaling [44], provides a simple post-processing approach to adjust the softmax probability scores of the trained models. Other lines of methods have proposed training objectives to enforce the model to produce less confident scores, either in the predictions space [45, 43], logits [32, 31], or modifying the ground truth labels [34, 35].
@@
Conformal Prediction (CP) [60], an alternative strategy to quantify the uncertainty [49, 48, 1], is a statistical framework which offers several advantages over calibration methods. First, unlike most methods in calibration, CP works directly on the model predictions, offering an appealing solution for black-box models. Second, instead of simply modeling the correctness of the predicted probability, CP produces a set of predictions, including the most likely classes, which can be of much interest in certain problems. Last, CP methods have theoretical guarantees for the marginal coverage of the true class within the predicted set, under several assumptions, contrary to calibration approaches.
@@
Due to these properties, CP is gaining attention to conformalize the predictions of deep models [2, 1, 11, 50]. Nevertheless, albeit the efforts to study CP in large language foundation models [15], its impact on vision foundation models has been unexplored, besides the significant implications it may have on a variety of vision problems. The aim of this study is to shed light and provide important insights into this direction. To achieve this, we conducted an extensive empirical analysis of the performance of three common CP methods on eleven popular vision foundation models across multiple vision datasets. Our extensive experiments further explore common situations encountered in practice, assessing the impact on CP: under distributional drifts, after confidence calibration and in few-shot adaptation to novel downstream tasks. Our key observations are:
@@
    Across all the experiments, Adaptive Prediction Sets (APS) is the best CP approach in terms of empirical coverage, while Regularized Adaptive Prediction Sets (RAPS) presents the best alternative from a conformal set size standpoint.
@@
    Few-shot adaptation of vision-language models (VLMs) yields better conformalization than zero-shot predictions in ID data, with marginal gains on OOD.
@@ 2 Related Work
Foundation models for computer vision.  The landscape of foundation models has rapidly evolved in the last years. Traditionally, pre-trained convolutional networks based on ResNet architectures [18] were the main models used by the community. However, driven by the unprecedented advances in language models, e.g., GPT [7] or LLaMA [57], as well as the vast availability of image data online, there are groundbreaking advances in unimodal [8, 40, 26] and multimodal [46] foundation models for vision tasks, commonly based on vision transformers. These large pre-trained models aim to generalize across a broad span of visual tasks by pre-training on massive, diverse image datasets, exhibiting strong zero-shot and generalization capabilities to new tasks. For example, vision foundation models such as DINO [8, 40] rely on self-supervised learning strategies on large datasets, leading to excellent semantic understanding of visual content. On the other hand, CLIP [46] bridges the gap between language and vision modalities through contrastive learning, effectively allowing the model to understand images in the context of natural language prompts and enabling zero-shot capabilities.
@@
Quantifying the uncertainty of the predictions of deep networks has recently garnered considerable interest. From a calibration standpoint, popular strategies include post-hoc approaches [16, 12, 24], which map the logits or softmax predictions to smoother distributions, and explicit learning objectives [31, 32, 35, 34, 19, 37], which are integrated into the loss function. Nevertheless, a main limitation of calibration methods is that they lack theoretical guarantees of model performance. In contrast, CP has recently emerged as a promising alternative, which provides marginal coverage guarantees over unseen test samples [61, 51]. Specifically, CP resorts to a non-conformity score function (i.e., a measure of how “different” a particular data point is compared to a CP calibration dataset) to produce a finite prediction set, which is guaranteed to contain the true label with a user-specified confidence level. A central objective of the CP literature has been to improve either the set efficiency (i.e., smaller set sizes) or the class conditional coverage. For this purpose, several non-conformity scores have been presented [56, 48, 1, 49, 13, 11, 55], with [48, 1, 49] being popular methods widely studied. A straightforward solution directly uses the raw class softmax predictions to generate the prediction sets [49]. Adaptive Prediction Sets (APS) [48] provides an adaptive version, computing non-conformity scores by accumulating sorted softmax probabilities in descending order. To further improve the efficiency, RAPS [1] introduces an explicit regularization term, which penalizes non-conformity scores for unlikely classes.
@@
However, a main limitation of existing evaluations is the focus on more traditional models, usually trained on data collection that falls within the calibration and test data points distribution. Despite this transfer learning framework not necessarily affecting the marginal guarantees provided by CP, how it affects its efficiency and conditional coverage remains to be explored. Thus, quantifying the uncertainty of their predictions is paramount given the rising popularity of foundation models in strategic domains. However, whereas uncertainty quantification from a calibration perspective has been scarcely studied [36, 66, 59], its exploration under CP is, to our knowledge, overlooked.
@@ 3.2 Non-conformity scores
Least Ambiguous Classifier (LAC) [49]  aims to construct the smallest possible set under the assumption that the output is correct. Intuitively, it can be interpreted as a thresholding of the output probabilities for each category. Thus, the non-conformity score can be constructed as:
@@
where $\rho(\mathbf{x},y)$ is the accumulated confidence of the categories more likely than the evaluated label $y$, i.e., $\rho(\mathbf{x},y)=\sum_{k^{\prime}\in\mathcal{Y^{\prime}}(\mathbf{x},y)}x_{k=k^{\prime}}$, with $\mathcal{Y^{\prime}}(\mathbf{x},y)=\{k|x_{k}>x_{k=y}\}$. Adaptive methods usually include $u\in\{0,1\}$, as a random variable to break ties to archieve exact marginal coverage.
@@
We refer the reader to the different works [1, 49, 48] for the respective conformal calibration coverage guarantees.
@@ 4.1 Experimental Setup
Models: We employ a total of 11 foundation models: two DINO [8] (DINO-S and DINO-B), four DINOv2 [40] (DINOv2-S, DINOv2-B, DINOv2-L, and DINOv2-G), three VICReg [3] (with ResNet-50, ResNet-50x2, and ResNet-200x2), and two VLMs (CLIP [46] and MetaCLIP [65]) based on ViT-B. Our main analysis is conducted on three popular vision datasets: CIFAR-10 [28], CIFAR-100 [28], and Imagenet [10], including its versions integrating domain shifts [21, 22, 62, 47]. Each dataset is split into two sets: one for training, and one for the conformal experiments. The latter is then split into one calibration set to tune the CP method, and one test set for evaluation. For few-shot, we adhere to the emerging CLIP few-shot literature [68, 53, 69], and evaluate models on 10 additional fine-grained and general concepts classification benchmarks: SUN397 [64], FGVCAircraft [33], EuroSAT [20], StanfordCars [27], Food101 [6], OxfordPets [41], Flowers102 [39], Caltech101 [14], DTD [9], and UCF101 [54].
Metrics: We resort to conformal set size (a.k.a. efficiency), minimum class-conditional coverage (MCCC), coverage gap (CovGap) and marginal coverage to benchmark the different methods. More details about metrics and implementation details can be found in Appendix A and Appendix B.
Figure 1: Relationship between the linear probing model accuracy and conformal set size (top) and the coverage gap (bottom) across different tasks of increasing complexity. From left to right: CIFAR-10, CIFAR-100 and ImageNet.
Adaptation to target tasks.  The foundation models used in this study have been pre-trained using different strategies. Nevertheless, they need to be adapted for novel tasks, as their pre-trained versions do not accommodate classification tasks, i.e., there is no classification head. To do this, foundation models are frozen, and a linear probing (LP) head (one linear layer followed by a softmax activation function) is trained on each dataset by optimizing a cross-entropy loss (more details in Appendix B).
@@ 4.2 Results
To gain insights into the factors influencing the efficacy of CP in vision foundation models, we design four experiments. First, we explore the impact of CP in standard scenarios, where a large calibration set is available to conform to the predictions of different models. Then, we challenge the status quo of CP and alter the conformal sets to accommodate real-world scenarios by including domain shifts. Furthermore, since confidence calibration is significantly linked to CP, we explore the impact of model calibration on CP performance. Last, we examine CP when adapting a very popular VLM, i.e., CLIP, to novel tasks.
@@ 4.2.1 Performance in the General Setting
First, we study the performance of 11 vision foundation models paired with CP methods under the standard setting, and on the three datasets, which presents ideal conditions: a sufficiently large calibration set, and absence of distributional drifts between calibration and test sets. We aim to determine whether we can prescribe a winner solution in this scenario and which factors can help identifying it.
@@
Fig. 1 depicts the relationship between the linear probing performance of each model and the conformal set size (top), and coverage gap (bottom). The initial observation highlights a clear trend: higher-performance models tend to produce smaller prediction sets, regardless of the conformal method used. Nevertheless, while set efficiency (i.e. size) is typically considered as a sufficient condition in most prior literature in conformal prediction [1], our analysis revealed that the relation with the accuracy does not consistently hold when examining other metrics. In particular, Fig. 1 (bottom) exposes that different CP methods yield mixed results for the coverage gap, which do not correlate directly with accuracy across all CP approaches. While APS appears to be almost unaffected by the model performance, RAPS clearly benefits from a strong performance, and LAC is negatively affected by more accurate models as the dataset becomes more complex.
@@
Regarding comparison between the CP methods, if we consider the set size, APS is clearly outperformed by the other approaches, particularly on ImageNet, whereas LAC provides the smallest prediction sizes, closely followed by RAPS. Indeed, RAPS is specifically designed to reduce the conformal set size of APS. However, the coverage gap results indicate that this comes at the cost of increasing the range of the class-conditional coverage. Below, we analyze the underlying causes that may explain this behaviour.
@@
RAPS class-conditional coverage, and therefore coverage gap, are more sensitive to the model’s accuracy. Let us assume we have two models, $M_{1}$ and $M_{2}$, in a multi-class classification problem, whose accuracies are $Acc_{1}$ and $Acc_{2}$, respectively. For each class $y\in\mathcal{Y}$, we refer to $C_{M_{i}}(y)=\mathbb{P}(Y=y|Y\in\mathcal{S}_{M_{i}}(X))$ as the class-conditional coverage for $y$ under model $M_{i}$, which measures the probability that predictions include the true label when the true label is $y$. Furthermore, let $\delta_{M_{1}}=\min_{y\in\mathcal{Y}}C_{M_{1}}(y)$ and $\delta_{M_{2}}=\min_{y\in\mathcal{Y}}C_{M_{2}}(y)$ denote the minimum class-conditional coverage achieved by each model. Under this scenario, we argue that due to the penalty in RAPS, if $Acc_{1}<Acc_{2}$, then $\delta_{M_{1}}<\delta_{M_{2}}$ when using RAPS as a conformal prediction method. In particular, $M_{1}$ (with lower accuracy) needs to expand its prediction sets for certain classes to meet the marginal coverage target 1-$\alpha$. However, the penalty term encourages small prediction sets, limiting an excessive number of classes. Thus, for some difficult classes, model $M_{1}$ may still potentially fail to meet the target coverage, as the enforced penalty discourages overly large sets. This ultimately results in lower coverage rates for those specific classes than $M_{2}$. In contrast, since APS does not include any regularization term that encourages small set sizes, it will compensate for lower performing models by increasing its set sizes, ultimately attaining higher class-conditional coverages.
@@
To confirm this hypothesis, we perform the following experiment. First, we identify the class with the lowest minimum class-conditional coverage obtained by RAPS for each model $m$, which we refer to as $y^{R}_{m}$, and find its corresponding set size (both represented blue circles in Fig. 2). Then, we identify the class-conditional coverage and set size of the class $y^{R}_{m}$ provided by APS across all models, whose values are shown as pink circles in both plots of Fig. 2. Last, we also include the minimum class-conditional coverage obtained by APS across each method, depicted with a cross. Note that the minimum class-conditional coverage from APS does not necessarily correspond to the coverage of class $y^{R}_{m}$. Upon close examination of these results, we observe that, indeed, for models presenting lower accuracy (those with smaller circles), the gap between the class-conditional coverage for the worst class in RAPS and the same class in APS is consistently larger than for more accurate models (larger circles). This empirical evidence shows that the minimum class-conditional coverage is substantially reduced because RAPS is constrained from expanding its predictions set size. This effect is more pronounced in less accurate models, where the true class may rank far away from the maximum allowable set size in the softmax predictions. In contrast, APS tolerates worse models by increasing the set size, which ultimately degrades the set efficiency but yields better class-conditional coverages.
Last, LAC presents structural differences with RAPS and APS, as it lacks an adaptive mechanism, relying on a uniform fixed threshold. Thus, LAC may yield inconsistent coverage rates across classes, resulting in high variability in the class-conditional coverage and thus in the coverage gap.
@@
$\text{ViT}_{\text{DINO-S}}$ 74.92 3.27 10.02 4.23 0.433 0.477 0.412
$\text{ViT}_{\text{MetaCLIP}}$ 75.80 2.39 9.43 2.84 0.479 0.535 0.467
@@
Table 1: SSL vs supervised learning. Results on ImageNet obtained by CLIP (ViT-B), MetaCLIP (ViT-B), and DINO-S (ViT-S) and a ViT-B trained in a supervised manner on ImageNet.
Following this analysis, we are also interested in determining whether a network pre-trained following a more traditional approach (i.e., standard supervised fine-tuning) offers similar conformal capabilities to self-supervised and contrastive ones. In particular, we select a ViT-B pre-trained on ImageNet, which is the same architecture as the visual encoder of the different foundation models. The results from this analysis (Table 1) reveal that, despite obtaining lower classification accuracy when using LP on the different foundation models, CP methods typically yield better performance than in $\text{ViT}_{\text{ImageNet}}$. These differences are significant under the APS approach, where the set size is significantly degraded on $\text{ViT}_{\text{ImageNet}}$. Moreover, the class-conditional coverage is also substantially affected, with nearly 6% decrease compared to the best model.
@@
Figure 4: Evaluation under domain-shift. Set size($\downarrow$), coverage($\uparrow$), and MCCC($\uparrow$) across three CP methods and three foundation models. ImageNet versions are sorted based on OOD performance in [53].
To further delve into these differences, we compute, for each test sample, the difference between the conformal set size for APS when applied to $\text{ViT}_{\text{ImageNet}}$ and $\text{ViT}_{\text{CLIP}}$ models, whose distribution is depicted in Fig. 3 (additional results in Appendix E). These values confirm that set size differences are not derived from a small set of isolated outliers but from a considerably large group of samples that see their conformal set increase when using the ViT trained in a supervised manner. These results suggest that the strategies used to train foundation models yield better CP metrics, resulting in conformalized models that can be deployed more safely on critical scenarios. It is important to stress that this study is limited due to the different dataset scales used for training (i.e., ImageNet alone is insufficient to train a foundation model). Our goal, however, is to understand the conformalization properties of readily available pre-trained models, regardless of how they were pre-trained.